AAVSOnet news, 2015-04-04

Sat, 04/04/2015 - 20:34

The following telescopes are performing nominally:  BSMs South, HQ, NM, Hamren, Berry; Coker30, OC61.

SRO has had dome problems recently.  Either the shutter control malfunctions, or the dome rotation malfunctions.  In most cases, the resident astronomer (Mike Shade) or John Gross can get things operational for the current night, but we're diagnosing the problems and should have a smoothly operating system soon.  If your images seem vignetted, this may be the reaon.

Coker30, while listed as nominal above, actually has had some recent tracking problems.  Bill Goff is aware of the situation and is working on it.  His water pump also recently failed, so we can't go quite as cold as before, but I'm sending him a new pump.

TMO61's problems continue.  Jon Holtzman can get things running as long as he doesn't communicate with the JMI focuser (which means no autofocusing).  There are a couple of workarounds that he is currently exploring.

BSM Argentina has been waiting for a replacement hand-controller for about a year.  We're hoping to get this system running again soon.

We are resuming testing of the eShel spectrograph on OC61 this week.  UC graduate student Byron Engler is heading to Mt. John as I write to work on the spectrograph.  This will likely mean a disruption of OC61 programs for the next week as the system gets reconfigured.  I expect the eShel to get used heavily during bright time.

Matt has been working with Ido Bareket on Bareket30.  I hope to have this system added to AAVSOnet shortly.

The diffraction gratings at BSM_South and BSM_Berry have been used extensively for monitoring of the recent southern novae.  They are a great way of obtaining spectral resolution on bright objects, and can be requested for your observing program.  We also have gratings on Coker30, BSM_NM and TMO61.  We hope to have the LISA spectrograph (donated by Bart Staels and Valère Van Ruymbeke) available later this year on TMO61.

Bill Stein is going to be sending us the Halpha filter from BSM_NM shortly.  Ulisse Munari has written a paper suggesting that the optimal monitoring of a nova can be done with Stromgren y, Halpha and O[III].  I am sending him our Halpha and O[III] filters, and Don Goldman has sent me a Stromgren y filter for testing.  After I get the spectral scans of the 3 filters from Ulisse, they will be installed at BSM_NM for use.  We currently have a half-dozen bright novae that can be followed, and they will take precedence with these 3 filters.

Gary Walker is working with Mike Linnolt to automate the BSM_Hamren enclosure.  This should give us about twice as many useful hours with that system.  Gary also spent several hours with the BSM_HQ enclosure, improving the weather tightness and the open/close functionality.

I've been talking with Dipankar Maitra at Wheaton College, and it looks like we may gain access to Grove Creek Observatory in Australia.  This will give us larger aperture access, along with OC61, in the southern hemisphere.

If anyone has problems with getting their AAVSOnet data, PLEASE contact me or post on this forum.  This is a great network, and I want it to continue to improve.  A few complaints help to highlight areas for improvement!